OpenStudy (anonymous):

write down the first fou terms in the expansion of (1+2x)^8. By substituting an appropriate value of x (which should be stated), find an approximate value of 1.02^8. State the degree of accuracy of your answer

OpenStudy (anonymous):

I got 1 + 16x + 112x^2 + 448x^3,, but I don't know how to find the approximation any ideas??? :)

OpenStudy (anonymous):

if x = 0.01, \(1+2x=1+2(0.01)=1+0.02=1.02\)

OpenStudy (anonymous):

so, in the expansion, plug in x=0.01 and find the approximation